### v1

@@ Abstract
This paper studies the problem of CPRP, concept prerequisite relation prediction, which is a fundamental task in using AI for education. CPRP is usually formulated into a link-prediction task on a relationship graph of concepts and solved by training the graph neural network (GNN) model. However, current directed GNNs fail to manage graph isomorphism which refers to the invariance of non-isomorphic graphs, reducing the expressivity of resulting representations. We present a permutation-equivariant directed GNN model by introducing the Weisfeiler-Lehman test into directed GNN learning. Our method is then used for CPRP and evaluated on three public datasets. The experimental results show that our model delivers better prediction performance than the state-of-the-art methods.
@@
Concept Prerequisite Relation, permutation-equivariant GNNs, Weisfeiler-Lehman Test, Directed Graph Learning, AI for Education.
@@ 1 Introduction
There are numerous approaches to solving the link prediction problem, including probabilistic models, spectral clustering, evolutionary algorithms, and deep graph learning models Kumar et al. (2020). Currently, graph neural networks (GNNs) have become the benchmark method and presented state-of-the-art performance in link prediction Cai et al. (2021), including many GNN-based solutions to CPRP Roy et al. (2019); Jia et al. (2021); Sun et al. (2022); Mazumder et al. (2023). To improve the capability of GNNs, Long et al. (2022) proposed to pre-train the node features via the method of graph reconstruction, achieving improved performance on two biological prediction tasks and effectively reducing training costs. Chamberlain et al. (2022) proposed to utilize subgraph sketches to pass messages in subgraph GNNs, delivering higher accuracy and lower computation costs in line prediction tasks. This model mitigated the expressive limitations, such as the inability to count triangles and distinguish automorphic nodes. Persistent homology was also adopted in the work of Yan et al. (2021) to extract topological information from graphs, which was integrated with node features to enhance the expressive power of GNNs for link prediction. Besides, the Weisfeiler-Lehman test was introduced by Morris et al. (2019) to improve the capability of differentiating graph isomorphism in undirected GNN learning Huang et al. (2022).
@@
In this paper, we extended the Weisfeiler-Lehman test-based GNN model, i.e., SpeqNets recently introduced by Morris et al. (2022), into the directed graphs for CPRP. The proposed framework contributes to learning the permutation-equivariant direct GNNs and improving the prediction performance of CPRP by distinguishing non-isomorphism graphs. Experimental results on three public datasets manifest that our method performs better than the state-of-the-art methods Sun et al. (2022); Jia et al. (2021); Roy et al. (2019).
@@ 3.2.1 KC Graph Construction
3.2.2 Weisfeiler Leman Guided Dircted GNNs

### v2

@@ Abstract
This paper studies the problem of CPRP, concept prerequisite relation prediction, which is a fundamental task in using AI for education. CPRP is usually formulated into a link-prediction task on a relationship graph of concepts and solved by training the graph neural network (GNN) model. However, current directed GNNs fail to manage graph isomorphism which refers to the invariance of non-isomorphic graphs, reducing the expressivity of resulting representations. We present a permutation-equivariant directed GNN model by introducing the Weisfeiler-Leman test into directed GNN learning. Our method is then used for CPRP and evaluated on three public datasets. The experimental results show that our model delivers better prediction performance than the state-of-the-art methods.
@@
Concept Prerequisite Relation, permutation-equivariant GNNs, Weisfeiler-Leman Test, Directed Graph Learning, AI for Education.
@@ 1 Introduction
There are numerous approaches to solving the link prediction problem, including probabilistic models, spectral clustering, evolutionary algorithms, and deep graph learning models Kumar et al. (2020). Currently, graph neural networks (GNNs) have become the benchmark method and presented state-of-the-art performance in link prediction Cai et al. (2021), including many GNN-based solutions to CPRP Roy et al. (2019); Jia et al. (2021); Sun et al. (2022); Mazumder et al. (2023). To improve the capability of GNNs, Long et al. (2022) and Chamberlain et al. (2022) proposed to pre-training node features through the method of graph reconstruction and utilizing subgraph sketches to pass messages in subgraph GNNs, aiming to enhance the capabilities of GNNs and reduce computational costs. Thess models mitigated the expressive limitations, such as the inability to count triangles and distinguish automorphic nodes. Persistent homology was also adopted in the work of Yan et al. (2021) to extract topological information from graphs, which was integrated with node features to enhance the expressive power of GNNs for link prediction. Besides, the Weisfeiler-Leman test was introduced by Morris et al. (2019) to improve the capability of differentiating graph isomorphism in undirected GNN learning Huang et al. (2022).
@@
In this paper, we extended the Weisfeiler-Leman test-based GNN model, i.e., SpeqNets recently introduced by Morris et al. (2022), into the directed graphs for CPRP. The proposed framework contributes to learning the permutation-equivariant direct GNNs and improving the prediction performance of CPRP by distinguishing non-isomorphism graphs. Experimental results on three public datasets manifest that our method performs better than the state-of-the-art methods Sun et al. (2022); Jia et al. (2021); Roy et al. (2019).
@@ 3.2.1 KC Graph Construction
3.2.2 Weisfeiler Leman Guided Directed GNNs